OpenStudy (albert0898):

If cot theta = 2/3 and csc theta > 0 , what is the numerical value of sec theta?

OpenStudy (mathstudent55):

\(\cot \theta = \dfrac{adj}{opp}\) \(\sec \theta = \dfrac{hyp}{adj} \)